Quizzes are worth 15%
Exams are worth 55%
Projects are wrorth 25%
Attendance is 5%
On quizzes I got 126 out of 150 possible points
On three Exams I got 79% 83%, 93% Out of 100%
On Proects I got 25 out of 27 possible points
I have perfect attendance
What is my grade in the class?

Answer by math_tutor2020(3817)


On the quizzes you got 126 out of 150, so that leads to 126/150 = 0.84 = 84%
Quizzes are worth 15% of your grade, so you have 0.84*0.15 = 0.126 = 12.6% of your grade so far just with the quizzes only.

Exams are worth 55%
Let's say each of the three exams has the same weight. This would mean each exam is worth (55%)/3 = 18.33% approximately
Apply this new weight to each exam score and we get:
0.79*0.1833 = 0.144807
0.83*0.1833 = 0.152139
0.93*0.1833 = 0.170469
Those add up to
0.144807+0.152139+0.170469 = 0.467415


So roughly 46.7415% of your grade is from the exams

Adding that to the quiz percentage we got, we update the total to
12.6% + 46.7415% = 59.3415%

Now onto the projects
They are worth 25%
You got 25/27 = 0.925926 approx
So, 0.25*0.925926 = 0.2314815
This means roughly 23.14815% of your grade is from the projects completed.

Updating the total gets us
59.3415% + 23.14815% = 82.48965%

Finally, you got perfect attendence. So that means you scored 100% on the attendence record and the percent contributed is 1*0.05 = 0.05 = 5%
So we add on 5% to get the final grade being
82.48965% + 5% = 87.48965%

That rounds to an 87%
------------------------------

Here's how you can do the calculation in one single expression:

0.15*(126/150)+(0.55/3)*(0.79+0.83+0.93)+0.25*(25/27)+0.05*1 = 0.87498148148149
This converts to 87.498148148149%

The result here is slightly off from the previous result of 87.48965% due to rounding error. Both results however round to 87% when rounding to the nearest whole percentage.

------------------------------

Answer: Final Grade = 87%